# VideoJAM: Joint Appearance-Motion Representations for Enhanced Motion Generation in Video Models

### v1

@@ 2 Related Work
Diffusion models (Ho et al., 2020) revolutionized visual content generation. Beginning with image generation (Dhariwal & Nichol, 2021; Rombach et al., 2022; Ho et al., 2022a; Black Forest Labs, 2024; Dai et al., 2023; OpenAI, 2024), editing and personalization (Gal et al., 2022; Ruiz et al., 2023; Chefer et al., 2024b; Sheynin et al., 2024; Singer et al., 2024; Chefer et al., 2024a), and more recently video generation. The first efforts to employ diffusion models for videos relied on model cascades (Ho et al., 2022b; Singer et al., 2023) or direct “inflation” of image models using temporal layers (Guo et al., 2023; BarTal et al., 2024; Wu et al., 2023). Other works focused on adding an auto-encoder for efficiency (Blattmann et al., 2023b; An et al., 2023; Wang et al., 2023), or conditioning the generation on images (Blattmann et al., 2023a; Zhang et al., 2023; Xing et al., 2023; Girdhar et al., 2024; Hong et al., 2022). Recently, the UNet backbone was replaced by a Transformer (Polyak et al., 2024; Brooks et al., 2024; Genmo, 2024; Gupta et al., 2023; HaCohen et al., 2024), mostly following Diffusion Transformers (DiTs) (Peebles & Xie, 2023).
@@
The gap between pixel quality and temporal coherence is a prominent issue (Ruan et al., 2024; Brooks et al., 2024; Liu et al., 2024b; Kang et al., 2024). Previous works explored motion representations to improve video generation in different contexts. Some methods use them as input for guidance or editing (Geng et al., 2024; Ma et al., 2023; Liu et al., 2024a; Cong et al., 2023). Note that their objective differs from ours since we aim to teach models a temporal prior rather than taking it as input. Other methods increase the amount of motion by separating content and motion generation (Ruan et al., 2024; Qing et al., 2023). Finally, most similar to our approach, recent works use motion representations to improve motion coherence in image-to-video generation (Shi et al., 2024; Wang et al., 2024), but these are limited to models conditioned on images.
@@ 4.3 Inner-Guidance
Following Ho & Salimans (2022), we jointly train the model to be conditional and unconditional on both auxiliary signals, $y,d$ by randomly dropping out the text in $30\%$ of the training steps, and the optical flow in $20\%$ of the steps (setting $d=\textbf{0}$), to facilitate the guidance formulation during inference,
@@ 5 Experiments
We conduct qualitative and quantitative experiments to demonstrate the effectiveness of VideoJAM. We benchmark our models against their base (pre-trained) versions, as well as leading proprietary and open-source video models, to highlight the enhanced motion coherence achieved by our framework.
@@
Baselines  We consider a wide variety of state-of-the-art models, both proprietary and open-source. In the smaller category, we include CogVideo2B, CogVideo5B (Hong et al., 2022), PyramidFlow (Jin et al., 2024), and the base model DiT-4B. In the larger category, we evaluate leading open-source models (Mochi (Genmo, 2024), CogVideo5B) and proprietary models with external APIs (Sora (Brooks et al., 2024), Kling 1.5 (KlingAI, 2024), RunWay Gen3 (RunwayML, 2024)), along with the base model DiT-30B [^1] The leading baselines were selected using the video leadboard.
Qualitative experiments  Figures 159 present results obtained using VideoJAM-30B. The results demonstrate a wide variety of motion types that challenge existing models such as gymnastics (e.g., air splits, jumps), prompts that require physics understanding (e.g., fingers pressed into slime, basketball landing in a net), etc.
@@
Quantitative experiments  We evaluate appearance and motion quality, as well as prompt fidelity using both automatic metrics and human evaluations. In all our comparisons, each model runs once with the same random seed for all the benchmark prompts. For the automatic metrics, we use VBench (Huang et al., 2024), which assesses video generators across disentangled axes. We aggregate the scores into two categories- appearance and motion, following the paper. The metrics evaluate the per-frame quality, aesthetics, subject consistency, the amount of generated motion, and motion coherence. More details on the metrics and their aggregation can be found in App. C.1.
@@
+VideoJAM-4B - - - 74.9 93.7
Ablations  We ablate the primary design choices of our framework. First, we ablate the use of text guidance and motion guidance in our inner guidance formulation (by setting $w_{2}=0$, $w_{1}=0$ in Eq. 8, respectively). Next, we ablate the use of motion prediction during inference altogether, by dropping the optical flow at each inference step ($d=\textbf{0}$). Finally, we ablate our guidance formulation by replacing it with the InstructPix2Pix (IP2P) guidance (Brooks et al., 2023) (see Sec. 2, App. A). Note that the results of the DiT models in Tabs. 12 also function as ablations, as they ablate the use of VideoJAM during training and inference.
@@
Limitations  While VideoJAM significantly improves temporal coherence, challenges remain (see Fig. 7). First, due to computational constraints, we rely on both limited training resolution and RGB motion representation, which hinder the model’s ability to capture motion in “zoomed-out” scenarios where moving objects occupy a small portion of the frame. In these cases, the relative motion magnitude is reduced, making the representation less informative (Eq. 5). For example, in Fig. 7(a), no parachute is deployed, and the motion appears incoherent. Second, while motion and physics are intertwined, leading to improved physics, our motion representation lacks explicit physics encoding. This limits the model’s ability to handle complex physics of object interactions. For example, in Fig. 7(b), the player’s foot does not touch the ball before it changes trajectory.
@@ 6 Conclusions
$\nabla\log p_{\theta}(x|c_{1},c_{2})=\nabla\log p_{\theta}(c_{1}|c_{2},x)+\nabla\log p_{\theta}(c_{2}|x)p_{\theta}(x)-\nabla\log p_{\theta}(c_{1},c_{2}),$ (11)
@@ Appendix F VideoJAM-bench Prompts
    A woman drinks from a water bottle in a forest. The woman has fair skin and brown hair. She is wearing a black jacket and black and white gloves.

### v2

@@ 2 Related Work
Diffusion models (Ho et al., 2020) revolutionized visual content generation. Beginning with image generation (Dhariwal & Nichol, 2021; Rombach et al., 2022; Ho et al., 2022a; Black Forest Labs, 2024; Dai et al., 2023; OpenAI, 2024), editing and personalization (Gal et al., 2022; Ruiz et al., 2023; Chefer et al., 2024b, a), and more recently video generation. The first efforts to employ diffusion models for videos relied on model cascades (Ho et al., 2022b; Singer et al., 2023) or direct “inflation” of image models using temporal layers (Guo et al., 2023; BarTal et al., 2024; Wu et al., 2023). Other works focused on adding an auto-encoder for efficiency (Blattmann et al., 2023b; An et al., 2023; Wang et al., 2023), or conditioning the generation on images (Blattmann et al., 2023a; Girdhar et al., 2024; Hong et al., 2022). Recently, the UNet backbone was replaced by a Transformer (Polyak et al., 2024; Brooks et al., 2024; Genmo, 2024; HaCohen et al., 2024), mostly following Diffusion Transformers (DiTs) (Peebles & Xie, 2023).
@@
The gap between pixel quality and temporal coherence is a prominent issue (Ruan et al., 2024; Brooks et al., 2024; Liu et al., 2024b; Kang et al., 2024). Previous works explored motion or physics-based signals to improve video generation. Some methods use them as input for guidance or editing (Geng et al., 2024; Ma et al., 2023; Liu et al., 2024a; Cong et al., 2023; Montanaro et al., 2024). Note that their objective differs from ours since we aim to teach models a temporal prior rather than taking it as input. Other methods suggest separating content and motion generation (Tulyakov et al., 2018; Ruan et al., 2024; Qing et al., 2023; Shen et al., 2024; Jin et al., 2024b). In the context of image generation, Liu et al. (2023) improves the realism of human images by incorporating spatial priors such as surface-normal and depth maps in addition to a conditioning skeleton given as input. Finally, most similar to our approach, recent works use motion representations to improve coherence in image-to-video generation (Shi et al., 2024; Wang et al., 2024), but these are limited to models conditioned on images.
@@ 4.3 Inner-Guidance
Following Ho & Salimans (2022), we jointly train the model to be conditional and unconditional on both auxiliary signals, $y,d$ by randomly dropping out the text in $30\%$ of the training steps, and the optical flow in $20\%$ of the steps (setting $d=\textbf{0}$). In particular, when the optical flow signal is dropped, the loss is derived with respect to the appearance only, resulting in the training scheme described in Sect. 4.1. Thus, the overall guidance formulation during inference becomes,
@@ 5 Experiments
We conduct qualitative and quantitative experiments to demonstrate the effectiveness of VideoJAM. We benchmark our models against their base (pre-trained) versions, as well as leading proprietary and open-source video models, to highlight VideoJAM’s enhanced motion coherence.
@@
Baselines  We consider a wide variety of state-of-the-art models, both proprietary and open-source. In the smaller category, we include CogVideo2B, CogVideo5B (Hong et al., 2022), PyramidFlow (Jin et al., 2024a), and the base model DiT-4B. In the larger category, we evaluate leading open-source models (Mochi (Genmo, 2024), CogVideo5B) and proprietary models with external APIs (Sora (Brooks et al., 2024), Kling 1.5 (KlingAI, 2024), RunWay Gen3 (RunwayML, 2024)), along with the base model DiT-30B. The leading baselines were selected using the video leaderboard.
5.1 Qualitative Experiments
Figures 159 present results obtained using VideoJAM-30B. The results demonstrate a wide variety of motion types that challenge existing models such as gymnastics (e.g., air splits), prompts that require physics understanding (e.g., fingers pressed into slime, basketball landing in a net), etc.
@@ 5.1 Qualitative Experiments
5.2 Quantitative Experiments
We evaluate appearance and motion quality, as well as prompt fidelity using both automatic metrics and human evaluations. In all our comparisons, each model runs once with the same random seed for all the benchmark prompts. For the automatic metrics, we use VBench (Huang et al., 2024), which assesses video generators across disentangled axes. We aggregate the scores into two categories- appearance and motion, following the paper. The metrics evaluate the per-frame quality, aesthetics, subject consistency, the amount of generated motion, and motion coherence. More details on the metrics and their aggregation can be found in App. C.1.
@@ 5.2 Quantitative Experiments
+VideoJAM-4B - - - 75.1 93.7
5.3 Ablation Study
We ablate the primary design choices of our framework. First, we ablate the use of text guidance and motion guidance in our inner guidance formulation (by setting $w_{2}=0$, $w_{1}=0$ in Eq. 8, respectively). Next, we ablate the use of motion prediction during inference altogether, by dropping the optical flow at each inference step ($d=\textbf{0}$). Finally, we ablate our guidance formulation by replacing it with the InstructPix2Pix (IP2P) guidance (Brooks et al., 2023) (see Sec. 2, App. A). Note that the results of the DiT models in Tabs. 12 also function as ablations, as they ablate the use of VideoJAM during training and inference.
@@ 5.3 Ablation Study
5.4 Limitations
While VideoJAM significantly improves temporal coherence, challenges remain (see Fig. 7). First, due to computational constraints, we rely on both limited training resolution and RGB motion representation, which hinder the model’s ability to capture motion in “zoomed-out” scenarios where moving objects occupy a small portion of the frame. In these cases, the relative motion magnitude is reduced, making the representation less informative (Eq. 5). For example, in Fig. 7(a), no parachute is deployed, and the motion appears incoherent. Second, while motion and physics are intertwined, leading to improved physics, our motion representation lacks explicit physics encoding. This limits the model’s ability to handle complex physics of object interactions. For example, in Fig. 7(b), the player’s foot does not touch the ball before it changes trajectory.
@@ 6 Conclusions
$\nabla\log p_{\theta}(x|c_{1},c_{2})=\nabla\log p_{\theta}(c_{1}|c_{2},x)+\nabla\log p_{\theta}(c_{2}|x)+\nabla\log p_{\theta}(x)-\nabla\log p_{\theta}(c_{1},c_{2}),$ (11)
@@ Appendix D VideoJAM-bench: Automatic Metrics Breakdown and Prompts
For completeness, Tab. 6 presents the VBench metrics not supported for external benchmarks, calculated on the VBench prompts. The reported results demonstrate that VideoJAM improves almost all aspects of generation, indicating once again that our framework does not sacrifice video quality for improved temporal coherence.
@@
Table 6: Breakdown of the remaining automatic metrics from VBench comparing our 30B model with its base model, DiT-30B, on VBench prompts. Our method improves almost all metrics, indicating that VideoJAM does not compromise appearance for motion enhancement.
Method Temporal Object Multiple Human Color Spatial Scene Appearance Temporal Overall
Flickering Class Objects Motion Relationship Style Style Consistency
DiT-30B 99.7 89.1 63.0 97.0 81.2 72.0 52.1 22.1 24.2 27.5
+VideoJAM-30B 99.7 90.7 70.0 99.0 91.1 73.3 49.7 23.3 24.4 27.6
@@ Appendix F VideoJAM-bench Prompts
    “A woman drinks from a water bottle in a forest. The woman has fair skin and brown hair. She is wearing a black jacket and black and white gloves.”